unicorn123:

What is the exact circumference of the circle? 10π ft 20π ft 40π ft 60π ft

unicorn123:

help

IhelpUuHELPme:

Can you please provide the image of the circle or atleast the info about it's radius or diameter?

unicorn123:

its A 20 feet the line in the cicle is staight

unicorn123:

https://static.k12.com/nextgen_media/assets/243627-VHS_PA_S2_03_L209_L309_Q1.gif clike this and its the cirlce

IhelpUuHELPme:

Alright so The formula for circumfrense is C = 2*pi*r or since we have the diameter, the formula is C = pi*d

IhelpUuHELPme:

pi is 3.14 your diameter is 20

IhelpUuHELPme:

so your equation would be C = 3.14*20

unicorn123:

thanks

IhelpUuHELPme:

What would you get when multiplying pi(3.14) with your diameter?

unicorn123:

6.3

IhelpUuHELPme:

6.3??

IhelpUuHELPme:

that's incorrect

IhelpUuHELPme:

so 3.14 *20.00 cross out the 0s at the end 3.14 times 2 is 614 without decimal add a 0 at the end from the 0 from 20 6140 the decimal point is two numbers from the right so your answer would be 61.4 accurately